OpenStudy (anonymous):

1. What is the value of (1/2)3? A. 1/8 B. 1/6 C. 11/2 D. 1/2 i know C. is Incorrect

OpenStudy (welshfella):

1/2 * 1/2 * 1/2

OpenStudy (welshfella):

- a half multiplied by itself 3 times

OpenStudy (anonymous):

so B. ???

OpenStudy (welshfella):

when multiplying fractions just multiply the top parts and multiply the bottom parts for example 1/3 8 1/3 = 1*1 / 3*3 = 1/9 its not B

OpenStudy (welshfella):

* that should be 1/3 * 1/3

OpenStudy (anonymous):

will it be 1/8 thier is no 1/9

OpenStudy (welshfella):

the 1/9 was just giving you another example how did you get 1/8?

OpenStudy (welshfella):

its 1 on the top of course because 1*1*1 = 1

OpenStudy (welshfella):

2*2*2 = ?

OpenStudy (anonymous):

2+2+2= 6

OpenStudy (welshfella):

NO - you are ADDING - you multiply them * means multiply

OpenStudy (anonymous):

oh :O

OpenStudy (welshfella):

(1/2)^3 means 1/2 to the power 3 which is 1/2 times 1/2 times 1/2

OpenStudy (anonymous):

so the answer is A.

OpenStudy (welshfella):

yes 1/8

OpenStudy (anonymous):

thx you

OpenStudy (welshfella):

yw